### v3

Abstract
We present a radiology-specific multimodal model for the task for generating radiological reports from chest X-rays. Our work builds on the idea that large language models can be equipped with multimodal capabilities through alignment with pre-trained vision encoders. On natural images, this has been shown to allow multimodal models to gain image understanding and description capabilities. Our proposed model (MAIRA-1) leverages a CXR-specific image encoder in conjunction with a fine-tuned LLM based on Vicuna-7B, and text-based data augmentation, to produce reports with state-of-the-art quality. In particular, MAIRA-1 significantly improves on the radiologist-aligned RadCliQ metric and across all lexical metrics considered. Manual review of model outputs demonstrates promising fluency and accuracy of generated reports while uncovering failure modes not captured by existing evaluation practices. More information and resources can be found on the project website: https://aka.ms/maira. ††*Corresponding authors: sthyland@microsoft.com, jaalvare@microsoft.com
1 Introduction
Large-scale pretraining of general-purpose image and language models has enabled the development of data-efficient large multimodal models. A common paradigm is to adapt a vision encoder to a pretrained LLM with varying levels of integration, as done by LLaVA (Liu et al., 2023b), InstructBLIP (Dai et al., 2023) and Flamingo (Alayrac et al., 2022). Here, we explore this paradigm for the specialised goal of generating radiology reports from images, namely generating the Findings section of a report based on a frontal chest X-ray and given the Indication for the study.
The Findings section contains the radiologist’s key observations about the image. To assist in their interpretation, a radiologist may refer to prior scans from the patient, other imaging modalities or views of the chest, or the patient’s clinical and medical history if available and not provided in the Indication. Hence, reporting on a chest X-ray requires synthesis of various data sources including multiple images taken over a period of time. For this study, we focus on the simplified single-image setting, acknowledging that the resulting model will be susceptible to generating ‘hallucinated’ references to prior studies (Ramesh et al., 2022; Bannur et al., 2023a) or lateral views.
Although the Findings section aims describe the image, the generation task crucially differs from image captioning. Chest X-ray reports must also establish the absence of findings, for example confirming that the insertion of a central venous line did not cause a pneumothorax. Secondly, the findings in a chest X-ray are not typical ‘objects’. Radiographic findings can be subtle variations in opacity against an otherwise-normal background of overlapping structures, requiring the extraction of fine-grained details from the image. Hence, findings generation remains a challenging multimodal task, requiring both the extraction of fine-grained details from the image and the generation of nuanced, radiology-specific language. A model which could generate the first draft of a radiology report has the potential to improve and expedite radiology reporting workflows (Huang et al., 2023), if it can first demonstrate a sufficient standard of clinical accuracy.
General-domain models demonstrably fail at the task of findings generation (Tu et al., 2023). We therefore develop a radiology-specific large multimodal model, which we call MAIRA-1. MAIRA-1 benefits from a pre-trained language model (Vicuna-7B (Chiang et al., 2023)), a radiology-specific image encoder (RAD-DINO (Pérez-García et al., 2024)), and the use of GPT-3.5 for data augmentation. By fine-tuning this model on a publicly available dataset of chest X-rays with corresponding reports (MIMIC-CXR (Johnson et al., 2019a, b)), we demonstrate the potential of large multimodal models for specialised radiology use-cases.
Concretely, we share the following outcomes:
1.
Performance competitive with existing state-of-the-art is possible without extremely large models or datasets.
2.
Design choices make a difference: keys to the success of this model are the use of a domain-specific image encoder with a larger image resolution and a more complex adapter layer, as well as the use of data augmentation and the Indications section of the report.
3.
Evaluation for this task remains challenging with heterogeneity across prior work. We report a wide variety of metrics to enable comparison, and demonstrate through stratified analyses the sensitivity of performance measures to test set characteristics.
We share examples of Findings sections generated by MAIRA-1 to highlight both its successes and limitations. By re-purposing pre-trained LLMs we have demonstrated what is possible in a constrained setting with a limited dataset. Relative to a radiologist, MAIRA-1 receives only a limited view of the patient, relying on a single chest X-ray and the Indication for the study. Hence, MAIRA-1 is a step towards realistic report-drafting systems, which will incorporate richer inputs such as previous images or other clinical information. We anticipate that with larger and cleaner training datasets, this flexible LLM-based approach can yield further gains.
2 Related work
Generalist foundation models
The paradigm of aligning image encoders to existing LLMs has achieved great success in the general domain. An example is LLaVA (Liu et al., 2023b), which combines a vision encoder with LLM using a simple adaptation layer applied to the image embedding, and employing visual instruction tuning to achieve general-purpose visual and language understanding. InstructBLIP (Dai et al., 2023) performs similar visual instruction tuning, but leverages a more powerful Q-former and a frozen LLM. Flamingo (Alayrac et al., 2022) couples the image and text more tightly, allowing a frozen LLM to cross attend to images. Flamingo additionally benefits from training on interleaved visual-text-video data. PaLM-E (Driess et al., 2023) is a multimodal decision making model, trained on text, image and sensor data, capable of embodied reasoning and decision making.
Medical/radiology adaptation of LLMs/MLLMs
There have been multiple efforts to specialise generalist foundation models to the medical domain or specifically for radiology applications. For example, Med-Flamingo (Moor et al., 2023) is based on OpenFlamingo (Awadalla et al., 2023; Alayrac et al., 2022) and was trained on images and captions from medical textbooks to perform few-shot visual question answering (VQA). Med-PaLM M (Tu et al., 2023) fine-tuned the proprietary PaLM-E model (Driess et al., 2023) on a broad collection of biomedical datasets, to build a system addressing diverse uni- and multimodal tasks over text, images, and genomics. LLaVA-Med (Li et al., 2023) proposed to adapt LLaVA (Liu et al., 2023b) with a curriculum of plain image–text pairs and generated multimodal instructions based on PubMed data. ELIXR (Xu et al., 2023) aligns a CXR encoder model, SupCon (Sellergren et al., 2022), with PaLM 2-S (Anil et al., 2023) through a multi-stage training process. The resulting model can perform tasks such as classification, semantic search, question answering and quality assurance. Radiology-GPT (Liu et al., 2023c) is a text-only model based on the Alpaca instruction-tuning framework (Taori et al., 2023), trained with radiology reports from the MIMIC-CXR dataset to perform findings-to-impression generation.
Radiology report generation
Given the long tail of possible observations in a chest X-ray, and the need for fine-grained description of findings, the generation of the narrative radiology report itself from the image is a promising target for machine learning systems (Wang et al., 2018). Such systems necessarily require a generative language modelling component; initially recurrent neural networks (Wang et al., 2018; Liu et al., 2019) giving way more recently to transformer architectures (Miura et al., 2021; Chen et al., 2020; Bannur et al., 2023a), including LLMs such as PaLM (Tu et al., 2023) and here Vicuna-7B (Chiang et al., 2023).
Given the need for clinical accuracy in the generated text, a line of work departs from a vanilla language-modelling loss and uses reinforcement learning (RL) to optimise for ‘clinically relevant’ rewards, based on the presence of specific findings (Liu et al., 2019; Irvin et al., 2019), or logical consistency (Miura et al., 2021; Delbrouck et al., 2022). A downside of such approaches is the reliance on models such as CheXbert (Smit et al., 2020) or RadGraph (Jain et al., 2021) to extract clinical entities, and the more complex optimisation problem. Here we demonstrate what is possible through plain auto-regressive language modelling alone, acknowledging that gains from more sophisticated training objectives or RL-based approaches are likely complementary.
Prior work has focused on generating different sections of the radiology report, sometimes the combination of both Findings and Impression (Wang et al., 2018; Chen et al., 2020; Jin et al., 2023; Yan et al., 2023), the Impression alone (Endo et al., 2021; Bannur et al., 2023a), or—as here—the Findings only (Miura et al., 2021; Delbrouck et al., 2022; Tanida et al., 2023; Nicolson et al., 2023; Tu et al., 2023). Jeong et al. (2023) and Yu et al. (2023) studied all three settings, providing evidence that the choice of section markedly impacts reported metrics, prohibiting comparison between variations of the task.
3 Method
In this section, we explain details of our task definition, dataset composition and preparation, modelling choices, and training and inference pipelines. We also describe our broad suite of evaluation metrics.
3.1 Task and data
Task description
We generate the main body section (Findings) of the report accompanying a chest X-ray. The Findings section contains the radiologist’s description of the normal and abnormal findings on the image. The image is typically accompanied by an Indication section providing the reason for the study, which may include some clinical history or a specific request from the referring clinician.
The most common chest X-ray view is a frontal image, either acquired from posterior to anterior (PA) or anterior to posterior (AP). Other imaging views of the chest are also routinely performed to aid in the diagnostic task, such as a lateral image. It is worth noting that, in isolation, the lateral view is not able to visualise the relevant anatomy with sufficient clarity to generate a comprehensive diagnostic report. Therefore, the frontal view remains the default view for the traditional chest X-ray clinical interpretation task(s) and our work similarly relies on the frontal view in line with most prior studies.111A notable exception is Tu et al. (2023), where the lateral view is used.
In addition, radiology reports typically contain a final Impression section summarising the actionable insights from the study, including suspected clinical diagnoses and recommendations for follow-up investigations. Such information cannot be fully gathered from the image alone—and often not even from the remainder of the report in isolation—relying heavily on the radiologist’s domain expertise, external data (e.g. patient history), and context of the study.
Hence, we focus on the task of generating specifically the Findings section of the report, given the Indication section (if available), and a single frontal chest X-ray image.
Dataset description
Train | Validation | Test | |
Subjects | 55,218 | 2,709 | 285 |
Studies | 131,613 | 6,471 | 2,210 |
DICOMs | 146,909 | 7,250 | 2,461 |
We train and evaluate on the MIMIC-CXR dataset (Johnson et al., 2019a, b), hosted on PhysioNet (Goldberger et al., 2000). This dataset from the Beth Israel Deaconess Medical Center in Boston, comprises a total of 377,110 DICOM images across 227,835 studies. Each imaging study is accompanied by a report. We process the DICOM images to remove all non-AP/PA scans. For each report, we extract the Findings and Indication sections, using the official MIMIC-CXR codebase.222https://github.com/MIT-LCP/mimic-cxr/blob/master/txt/section˙parser.py We discard all studies for which Findings could not be extracted, but we allow for missing Indication sections. We use the standard MIMIC-CXR test split in our experiments, however, we additionally exclude benchmark datasets MS-CXR (Boecking et al., 2022) and MS-CXR-T (Bannur et al., 2023b) from our training set. Table 1 describes the subject, study and DICOM counts in each split. To facilitate comparison, we provide the list of DICOM identifiers used in our test split as an ancillary file.
We derive our images from the original DICOM files, instead of the commonly used JPEG files from MIMIC-CXR-JPG (Johnson et al., 2019c), the latter of which may contain compression artefacts and further loss of detail from grayscale quantisation. Following the pre-processing pipeline from CLIP (Radford et al., 2021), images are resized to match the shortest edge to the input size of the image encoder, then the longest edge is centre-cropped to the same size. Intensities are normalised according to the image encoder’s training data.
Data augmentation using GPT
We use GPT-3.5 to paraphrase both the Findings and Indication sections of the training set as a data augmentation technique, resulting in an additional 131,558 reports.333We used a private, compliant deployment of OpenAI’s gpt-3.5-turbo version 0301 on Microsoft Azure (https://learn.microsoft.com/en-us/azure/ai-services/openai/overview), such that MIMIC data was not sent to public-facing servers. We instruct GPT-3.5 to rewrite the findings and indication while preserving the information and radiology style. Table 2 shows an example. Note that we leave the validation and test sets unchanged.
Original | INDICATION: _F with presyncope. r/o infection // ?pneumonia |
FINDINGS: AP and lateral chest radiograph demonstrates hyperinflated lungs. Cardiomediastinal and hilar contours are within normal limits. There is no pleural effusion or pneumothorax. No evidence of pulmonary edema. Lungs are without a focal opacity worrisome for pneumonia. There is no air under the right hemidiaphragm. | |
Paraphrased | INDICATION: Female patient with pre-syncope. Suspected infection, possibly pneumonia. |
FINDINGS: Chest radiographs show lungs with hyperinflation, but normal cardiomediastinal and hilar contours. No pleural effusion, pneumothorax, or pulmonary edema is observed. No focal opacity is present in the lungs that could indicate pneumonia. Additionally, there is no air under the right hemidiaphragm. |
3.2 Model architecture
Our model consists of an image encoder, a learnable adapter on top of the image features, and an LLM, following the LLaVA-1.5 architecture (Liu et al., 2023b, a). We use the radiology-specialised image encoder, RAD-DINO (Pérez-García et al., 2024). RAD-DINO is a ViT-B model (Dosovitskiy et al., 2020) with 87M parameters trained on 838k chest X-ray images. The input resolution is 518518 and the encoder patch size is 14. We employ Vicuna-7B (Chiang et al., 2023) as the LLM. We note from LLaVA-Med (Li et al., 2023) that this initialisation from an LLM pretrained only on language should lead to better performance than from one already trained on multimodal data. Our adapter is a multi-layer perceptron (MLP) with GELU activations (Hendrycks and Gimpel, 2016) and hidden size 1024 for all layers. The adapter weights are randomly initialised following the defaults from PyTorch v2.0.1.
We prompt the model with a system message and a human instruction interleaved with the relevant image. The model is trained to output the correct answer, i.e. the full Findings section of the report. We first embed the image into a sequence of image patch tokens using the image encoder, taking the embeddings from the last layer of the model and excluding the CLS token. The adapter MLP is applied to each embedding to transform these image tokens to the input space of the LLM. After tokenising and embedding the prompt and answer, we insert the image patch tokens at the specified location in the prompt (typically between the system message and human instruction). The instructions we use with MAIRA-1 are “{image placeholder} Provide a description of the findings in the radiology image given the following indication: {indication}” if we know the Indication for the image, otherwise “{image placeholder} Provide a description of the findings in the radiology image.”.
3.3 Training and inference
We train with a standard auto-regressive language modelling loss (cross-entropy) (Graves, 2013). We use similar hyperparameters to LLaVA-1.5 for training, i.e. tuning the LLM jointly with the randomly initialized adapter (Liu et al., 2023b). Unlike LLaVA-1.5 we do not have a precursor training step to pretrain the adapter (see Appendix A for an experimental comparison). We train for 3 epochs without any parameter-efficient fine-tuning techniques. We use a cosine learning rate scheduler with a warm-up of 0.03 and learning rate . The global batch size is 128. Based on the behaviour of validation metrics observed throughout training in our experiments, we take the final checkpoint for all runs. For inference, we decode in 32-bit precision up to 150 tokens.
3.4 Evaluation metrics
We evaluate the generated reports using both lexical and radiology-specific metrics, as described below. Radiology-specific metrics typically focus on how particular findings are described in the text, for example whether ‘consolidation’ is noted as present or absent. For this reason, they are more sensitive to clinically-relevant aspects of the generated report, rather than potentially-superficial variation in phrasing.
Lexical metrics
We employ a collection of traditional NLP metrics designed to quantify word overlap between generated and reference texts. In particular, ROUGE-L (Lin, 2004) quantifies the length of the longest common word subsequence relative to the lengths of predicted and reference reports. BLEU-4 (Papineni et al., 2002) is based on -gram precision (geometric mean for up to 4), with a brevity penalty to discourage too short predictions. Lastly, we report METEOR (Banerjee and Lavie, 2005), which first aligns individual words (unigrams) in the prediction and reference, while attempting to preserve their ordering; then computes a weighted harmonic mean of unigram precision and recall, with a penalty for fragmentation of consecutive word subsequences.444We run METEOR with the default parameters as originally presented by Banerjee and Lavie (2005).
CheXpert F1
This set of metrics uses the CheXbert automatic labeller (Smit et al., 2020) to extract ‘present’/‘absent’/‘uncertain’ labels for each of the 14 CheXpert pathological observations (Irvin et al., 2019) from a generated report and the corresponding reference. As done originally by Irvin et al. (2019), we compute two versions of this metric, mapping the ‘uncertain’ label to negative or positive, to enable comparison with papers that use either. The binary F1 score for each CheXpert category is then computed conventionally as the unweighted harmonic mean of precision and recall. We report the macro- and micro-averaged F1 scores over 5 major observations555Following Miura et al. (2021); Tu et al. (2023), and others, based on the CheXpert ‘competition tasks’ selected by Irvin et al. (2019), the subset of 5 major categories considered is: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. and all 14 observations, referring to them respectively as ‘[Macro/Micro]-F1-[5/14]’.
CheXbert vector similarity
This metric feeds the generated and reference reports through the CheXbert model (Smit et al., 2020), then calculates the cosine similarity between their embeddings (Yu et al., 2023). We compute this metric as well as RadGraph F1 and RadCliQ using the code released by Yu et al. (2023).666https://github.com/rajpurkarlab/CXR-Report-Metric/tree/v1.1.0
RadGraph-based metrics
The RadGraph model (Jain et al., 2021) parses radiology reports into graphs containing clinical entities (references to anatomy and observations) and relations between them. Introduced in Yu et al. (2023), the RadGraph F1 metric computes the overlap in entities and relations separately, then reports their average. Entities are considered to match if the text spans and assigned types are the same, and relations are matched if their endpoint entities and relation type are the same.
We also compute a variant F1 metric used in prior work (Delbrouck et al., 2022), to enable direct comparison. Specifically, we report their RGER score, which matches entities based on their text, type, and whether or not they have at least one relation. For this, we use the radgraph package.777https://pypi.org/project/radgraph/
RadCliQ
Also proposed by Yu et al. (2022), RadCliQ (Radiology Report Clinical Quality) is a composite metric that integrates RadGraph F1 and BLEU score in a linear regression model to predict the total number of errors that radiologists would identify in a report. In their study, this metric had the closest alignment with radiologists’ judgement of report quality. For consistency with prior work, we use version 0 of RadCliQ.
4 Experiments
In this section, we present an overview of the experimental design, results and ablations. The main experimental design is seen in Table 3 where we detail design choices related to data, pretraining, image encoders and adapters. To account for slight differences in the definitions of the test split in other studies (see Section 4.3), for all our experiments we report medians and 95% confidence intervals, estimated over 500 bootstrap samples of the MIMIC-CXR test set. For all metrics, higher values are better, with the exception of RadCliQ (indicated by ‘’ in the tables).
4.1 Adapting existing large multimodal models
MAIRA-1 training starts the multimodal alignment from scratch, fully relying on the findings generation task to align the image and text embeddings into a joint representation space. To measure the effect of fine-tuning after aligning on other data, we compare to baselines fine-tuning three existing large multimodal models: LLaVA-1.0 (Liu et al., 2023b) and LLaVA-1.5 (Liu et al., 2023a) in the general domain, and LLaVA-Med (Li et al., 2023) in the biomedical domain.
Table 4 (a) compares the performance of these settings. We note that while fine-tuning LLaVA-1.5 outperforms training a comparable model from scratch (Table 4, ‘CLIP+MLP-2’), this advantage is lost when we replace CLIP with a domain-specific image encoder (Table 4, ‘RAD-DINO+MLP-2’). Out of the different large multimodal models compared, we find that LLaVA-1.5 performs the best when fine-tuned for the findings generation task, although it lags behind MAIRA-1.
4.2 Which components are most beneficial?
MAIRA-1 differs from LLaVA-based multimodal models in its domain-specific image encoder, a deeper adapter module, and the use of a GPT-augmented dataset during training. Table 3 describes several experiments showing the additive effect of these optimizations included in MAIRA-1, with results presented in Table 4 (b). We start from an architecture mirroring LLaVA-1.5, using a pre-trained CLIP image encoder, a randomly initialized two-layer adapter, and Vicuna-7B as the LLM. We train the adapter and LLM jointly on in-domain radiology data to obtain an initial baseline (‘CLIP+MLP-2’). Note that the performance of this model is slightly worse than that of a model fine-tuned from LLaVA-1.5, due to the random initialisation of the adapter.
By switching the image encoder from CLIP to a domain-specific model (RAD-DINO) we observe an improvement across all metrics, more than compensating for the need to re-initialise the adapter. The use of the domain-specific encoder also increases the number of image tokens from 576 to 1369. By correspondingly increasing the size of the adapter from two to four layers (‘MLP-4’), we show further improvements across both clinical and lexical metrics.
As a final step, we augment the dataset with GPT-paraphrased samples (Section 3) to produce MAIRA-1. The effect of the GPT-augmentation is to increase clinical metrics while slightly harming lexical metrics. This could be explained by a mild distribution shift in the GPT-paraphrased reports respective to the original MIMIC reports. In Appendix B we demonstrate that gains from the use of GPT-paraphrased samples are not simply due to training longer.
4.3 How does MAIRA-1 compare to existing approaches?
Strict comparison with prior work is challenging due to variation in test set inclusion criteria and pre-processing steps, despite the existence of a ‘canonical’ test split for MIMIC-CXR. For example, Tu et al. (2023) starts from the official split, but includes lateral images paired with the study’s report as independent samples, resulting in a reported test size of 4,834 images. Yu et al. (2023) and Jeong et al. (2023) take only a single image for each study, resulting in 1,597 samples in the test set888We observed 2,210 studies after taking one image per study., and Tanida et al. (2023) follows the split provided by Chest ImaGenome (Wu et al., 2022). Recall that our test set size is 2,461 image-report samples. We consider a full reproducibility study in the style of Johnson et al. (2017) out of scope for this work, however to enable future comparison we share the image identifiers used in our test set in an ancillary file. Changes in the distribution of the test set can significantly impact reported numbers, as demonstrated in Section 4.4. We attempt to account for some of this variability by reporting 95% confidence intervals from bootstrap replicates on the test set when we compare with prior work, but numbers must be interpreted with caution.
Given the above caveats, in Table 5 we compare MAIRA-1 to prior work. For all lexical metrics, MAIRA-1 seemingly outperforms or matches prior state of the art (SOTA). The substantial increase in METEOR relative to Tanida et al. (2023) is notable. On clinical metrics, there is no single superior approach—Tu et al. (2023) report slightly superior scores on RadGraph-F1 while Delbrouck et al. (2022) substantially outperforms on RGER—this is perhaps expected given their model was optimised for RGER. Across the CheXbert-derived metrics, MAIRA-1 is comparable or superior across the fourteen-classes, but slightly underperforms on the five-class subset relative to Tu et al. (2023). We present class-stratified results for MAIRA-1 in Section 4.4 (Table 6). Promisingly across all clinical metrics, MAIRA-1 sets a new standard for the radiologist-aligned RadCliQ score.
Category | Metric | MAIRA-1 | SOTA | [ref.] | Param. count | Test set size |
Vision / LLM | ||||||
Lexical | ROUGE-L | 28.9 [28.4, 29.4] | (Tu et al., 2023) | 22B / 62B | 4,834 images | |
BLEU-1 | 39.2 [38.7, 39.8] | (Tu et al., 2023) | 22B / 62B | 4,834 images | ||
BLEU-4 | 14.2 [13.7, 14.7] | (Miura et al., 2021) | 8M / ? | 2,347 reports | ||
METEOR | 33.3 [32.8, 33.8] | (Tanida et al., 2023) | 26M / 355M | 32,711 images | ||
Clinical | RadGraph-F1 | 24.3 [23.7, 24.8] | (Tu et al., 2023) | 22B / 62B | 4,834 images | |
RGER | 29.6 [29.0, 30.2] | (Delbrouck et al., 2022) | 8M / 18M | 2,347 reports | ||
CheXbert vector | 44.0 [43.1, 44.9] | (Miura et al., 2021)* | 8M / ? | 1,597 images | ||
RadCliQ () | 3.10 [3.07, 3.14] | 3.277 | (Miura et al., 2021)† | 8M / ? | 1,597 images | |
Macro-F1-14 | 38.6 [37.1, 40.1] | (Tu et al., 2023) | 22B / 62B | 4,834 images | ||
Micro-F1-14 | 55.7 [54.7, 56.8] | (Tu et al., 2023) | 22B / 62B | 4,834 images | ||
Macro-F1-5 | 47.7 [45.6, 49.5] | (Tu et al., 2023) | 22B / 62B | 4,834 images | ||
Micro-F1-5 | 56.0 [54.5, 57.5] | (Tu et al., 2023) | 22B / 62B | 4,834 images | ||
Macro-F1-14+ | 42.3 [40.9, 43.6] | – | ||||
Micro-F1-14+ | 55.3 [54.3, 56.2] | – | ||||
Macro-F1-5+ | 51.7 [49.9, 53.1] | – | ||||
Micro-F1-5+ | 58.8 [57.4, 60.0] | (Tanida et al., 2023) | 26M / 355M | 32,711 images |
4.4 Stratified results
Performance depends on finding class
Table 6 shows a breakdown of MAIRA-1 performance based on the finding class. We use the standard 14 hierarchical classes initially proposed by Irvin et al. (2019) and used in the CheXbert labeller (Smit et al., 2020).
Looking at F1-score, we note that the Macro-F1-14 values we report elsewhere mask variance across classes. We see high-performing classes (Support Devices: 84.5, Pleural Effusion: 68.9, Cardiomegaly: 64.0) and indeed poorly performing classes (Enlarged Cardiomediastinum: 11.9, Pleural Other: 14.7, Pneumonia: 18.3). We note that these latter categories are rarer and more nebulously defined,999As acknowledged by Irvin et al. (2019), pneumonia is a clinical diagnosis which should strictly not be assessed from a chest X-ray alone. and may be subject to more noise in the output of the CheXbert labeller itself. For example, Cardiomegaly (37% prevalence) should imply an enlarged cardiomediastinum, and yet the latter category occurs in only 8% of cases.
Specificity and negative predictive value (NPV) are not measured in the Macro-F1-14 aggregate metric, where we see consistently high values. Compared to recall and precision (positive predictive value), we infer the model may under- or miss-call positive findings, but more reliably reports on the absence of findings. This is useful for findings such as pneumothorax (NPV 98.9) in an intensive care setting, where a chest X-ray may serve to confirm the absence of a pneumothorax after the insertion of a drain or tube.
Results differ between normal and abnormal studies
When there are no findings in a study, radiology reports are often formulaic, containing templated phrases such as “No [evidence of] acute cardiopulmonary process". This constitutes a simpler language generation task for the model, as reflected in the higher lexical metrics in Table 7. In addition, we observe higher clinical metrics in the no-finding subset, which we speculate could be related to the distribution shift between MIMIC-CXR dataset splits. As noted by Johnson et al. (2019c), studies with no findings are over-represented in the released training and validation splits, compared to the test set.
The model benefits from the indication section
Category | Metric | Has finding | No finding | Has indication | No indication |
% () | 78.3% (1928) | 21.7% (533) | 57.5% (1414) | 42.5% (1047) | |
Lexical | ROUGE-L | 27.6 [27.1, 28.1] | 33.4 [31.9, 34.9] | 32.7 [32.0, 33.4] | 23.6 [23.1, 24.1] |
BLEU-4 | 13.2 [12.7, 13.7] | 18.7 [17.4, 20.3] | 17.6 [16.9, 18.2] | 9.0 [8.6, 9.6] | |
METEOR | 31.9 [31.4, 32.5] | 38.5 [37.1, 40.0] | 36.8 [36.1, 37.6] | 28.6 [28.0, 29.3] | |
Clinical | RadGraph-F1 | 23.0 [22.4, 23.5] | 28.5 [27.0, 29.9] | 27.8 [27.0, 28.5] | 19.2 [18.6, 20.0] |
RGER | 28.5 [27.9, 29.0] | 33.7 [32.0, 35.3] | 33.5 [32.8, 34.3] | 24.3 [23.6, 25.0] | |
ChexBert vector | 42.3 [41.3, 43.2] | 50.3 [48.1, 52.2] | 47.3 [46.2, 48.3] | 39.5 [38.2, 40.8] | |
RadCliQ () | 3.19 [3.16, 3.22] | 2.79 [2.71, 2.88] | 2.88 [2.84, 2.92] | 3.41 [3.37, 3.45] | |
Macro-F1-14 | 38.1 [36.2, 39.9] | – | 39.1 [37.1, 41.5] | 36.9 [34.4, 39.2] | |
Micro-F1-14 | 57.0 [55.8, 57.9] | – | 56.3 [54.7, 57.8] | 55.0 [53.3, 56.6] | |
Macro-F1-5 | 48.8 [46.7, 50.8] | – | 46.8 [44.5, 50.0] | 47.3 [44.4, 50.4] | |
Micro-F1-5 | 57.5 [56.0, 59.0] | – | 56.6 [54.6, 58.6] | 55.5 [53.3, 57.8] |
The study indication is expected to substantially affect the contents of a radiology report. For example, as in Table 2, it may prompt the radiologist to report on specific types of abnormality that might not be routinely included, and is particularly relevant for deciding what to report as absent from the scan. MAIRA-1 uses the indication section of the report to help generate the findings section, whenever it is available with the study (66.3% in training and 57.5% in the test set). In Table 7 we show how performance varies in subsets of the test set with and without the indication section. Indeed, by leveraging the indication, MAIRA-1 is able to to generate radiology reports that are drastically more similar and more accurate with respect to the true report than in the subset without indication.
5 Examples
Figure 1 reproduces the example101010In this section we show the JPEG versions of images as released in MIMIC-CXR-JPG (Johnson et al., 2019c). However, as described in Section 3, MAIRA-1 ingests images derived from the original DICOMs. shown in Tu et al. (2023), comparing the output of MAIRA-1 with the best Med-PaLM M variant (84B) and the closest by model size (12B). For illustration, we also include the output of models not trained on the findings generation task (LLaVA-1.5 and LLaVA-Med).
Original report |
INDICATION: _-year-old male with intubation. No prior examinations for comparison. |
FINDINGS: Endotracheal tube terminates 6 cm from the carina. Nasogastric tube loops within the stomach, with side port within the gastric body. There are no significant pleural effusions or pneumothorax. Mild retrocardiac atelectasis is present. Heart size is normal. The aorta is tortuous. |
Model | Generated findings |
MAIRA-1 | Endotracheal tube terminates 5 cm above the carina. Nasogastric tube loops in the stomach, with side port near the GE junction. Heart size is normal. Aorta is tortuous and unfolded. There are no pleural effusions or pneumothorax. Lungs are clear. [No comment about atelectasis.] |
Med-PaLM M-12B | An endotracheal tube terminates 5 cm above the carina. There is no pneumothorax or large pleural effusion. Opacities in the medial right lower lobe and retrocardiac region may represent atelectasis aspiration or pneumonia. A calcified lymph node is noted in the aortopulmonary window. Cardiac size is normal. [No comment about NG/OG tube.] |
Med-PaLM M-84B | Endotracheal tube terminates 5 cm above the carina. Nasogastric tube courses below the diaphragm and inferiorly beyond the film. There is retrocardiac opacity which may represent atelectasis or aspiration. The heart size is normal. The aorta is tortuous and unfolded. No pleural effusions or pneumothorax. |
LLaVA-Med | The image is a chest X-ray taken on day 1 post-admission. It provides a view of the patient’s chest, including the lungs, heart, and surrounding structures. Chest X-rays are commonly used to help diagnose and monitor various conditions affecting the respiratory and cardiovascular systems, such as infections, fluid accumulation, or structural abnormalities. |
LLaVA-1.5 | The image displays a close-up view of a human body, focusing on the chest area. The chest is filled with various medical instruments, including a heart monitor and a defibrillator. The heart monitor is placed on the left side of the chest, while the defibrillator is positioned on the right side. In addition to the medical instruments, there are several wires and cords scattered throughout the scene. Some of these wires are attached to the medical devices, while others are likely part of the electrical connections for the instruments. The presence of these wires and cords suggests that the medical equipment is in use, possibly for monitoring or treating a patient. |
To conduct further error analysis, we sorted all examples in the test set by their RadCliQ score and manually reviewed the top and bottom cases with two consultant radiologists. The highest-scoring examples (lowest RadCliQ) were largely cases without findings, where the report follows a standard format; this is seen already in Table 7 when we stratify by whether the case has a finding. Figure 2 shows such an example; here MAIRA-1 has generated an almost-identical report to that of the original radiologist, with the exception of providing less detail on the sternal wires.
Original report | INDICATION: _-year-old man with cough and fever, question acute process. |
FINDINGS: The lungs are well expanded and clear. The cardiomediastinal silhouette, hilar contours, and pleural surfaces are normal. No pleural effusion or pneumothorax is present. Sternal wires are intact. | |
MAIRA-1 | FINDINGS: The lungs are well expanded and clear. The cardiomediastinal silhouette, hilar contours, and pleural surfaces are normal. No pleural effusion or pneumothorax is present. Median sternotomy wires and surgical clips are noted. |
A recurring but expected ‘failure’ mode we observe in MAIRA-1 is the generation of descriptions of change. This is a well-known problem in the generation of radiology reports (Bannur et al., 2023a; Ramesh et al., 2022): Standard clinical practice involves reporting using both current and prior studies precisely to enable assessment of change. This version of MAIRA-1 is trained on reports from single images, where reports often contain descriptions of changes compared to a previous study. This type of report is then also sometimes produced at test time, as illustrated in some of the following examples.
Figures 3 and 4 show success cases from MAIRA-1. In Figure 3, MAIRA-1 successfully avoids hallucinating a comparison to a prior study (‘Lung volumes are low.’), while Figure 4 shows examples of such hallucination (‘… pulmonary edema, new compared to the prior study.’), with other details largely correct.
Original report | INDICATION: _F with shortness of breath, congestive heart failure. |
FINDINGS: Lung volumes are low compared to the previous study. Left-sided AICD device is noted with single lead terminating in unchanged position in the right ventricle. Heart size appears at least mildly enlarged. The mediastinal and hilar contours are unremarkable. There is crowding of the bronchovascular structures without overt pulmonary edema. Streaky opacities in the lung bases likely reflect areas of atelectasis. No pleural effusion or pneumothorax is present. There are no acute osseous abnormalities. | |
MAIRA-1 | FINDINGS: Lung volumes are low. Left-sided AICD device is noted with single lead terminating in the region of the right ventricle. Heart size is mildly enlarged. Mediastinal and hilar contours are unremarkable. Crowding of bronchovascular structures is present without overt pulmonary edema. Patchy opacities in the lung bases likely reflect areas of atelectasis. No pleural effusion or pneumothorax is present. There are no acute osseous abnormalities. |
Original report | INDICATION: [No indication] |
FINDINGS: The patient is status post median sternotomy, CABG, and mitral valve replacement. The heart is mildly enlarged. The mediastinal contours are unchanged with calcification of the aortic knob again noted. Mild pulmonary edema appears progressed compared to the prior exam with small bilateral pleural effusions, also minimally increased compared to the prior exam. Left basilar opacification likely reflects atelectasis. There is no pneumothorax. No acute osseous abnormalities are identified. | |
MAIRA-1 | FINDINGS: The patient is status post median sternotomy and mitral valve replacement. Heart size is moderately enlarged but unchanged. The mediastinal contours are stable. There is mild interstitial pulmonary edema, new compared to the prior study. Small bilateral pleural effusions are also new. Retrocardiac opacity likely reflects atelectasis. There is no pneumothorax. No acute osseous abnormalities are visualized. |
Figure 5 shows a case with high (poor) RadCliQ score, however this does not reflect a failure of the model and rather a limitation of the evaluation. In this case, the original report contains negligible detail (‘Compared to the prior exam there is no significant interval change.’), whereas MAIRA-1 generates a full report which is largely correct. We note that examples of the converse also exist, where MAIRA-1 simply refers to the absence of change without further elaboration (for 491 studies, the Findings section is exactly the sentence ‘Compared to the prior study there is no significant interval change.’). This underscores a limitation of training with and evaluating on ‘noisy’ real-world datasets.
We further note that, in Figure 5, MAIRA-1 has generated a quantitative measurement (‘… tube is positioned 4cm above the carina’), which is not grounded in physical measurements of the image. This is also observed in Figure 1 from both MAIRA-1 and Med-PaLM M. Whereas a model may learn about an average field-of-view of the images seen during training, as well as certain correlations in the training reports, such measurements cannot be produced accurately without knowledge of physical and geometric parameters of the image acquisition.
Original report | INDICATION: [No indication] |
FINDINGS: Compared to the prior exam there is no significant interval change. | |
MAIRA-1 | FINDINGS: The endotracheal tube is positioned 4 cm above the carina, while the nasogastric tube is located below the diaphragm. The cardiac silhouette is enlarged with pulmonary vascular congestion. The left base shows opacification, indicating volume loss in the lower lobe and pleural effusion. The right base shows less prominent atelectatic changes and a possible small effusion. |
6 Discussion
We have presented results on MAIRA-1, a radiology-specialised large multimodal model designed for the task of generating the Findings section of a chest X-ray report.
Architecturally, MAIRA-1 consists of a frozen domain-specific image encoder based on a ViT-B (RAD-DINO), a four-layer feedforward adapter module, and the LLM Vicuna-7B. We train MAIRA-1 solely on the open-access MIMIC-CXR dataset, leveraging GPT-3.5 for text-based data augmentation. With these components, we demonstrate that performance competitive with existing state-of-the-art is possible, with either fewer parameters or a simpler training objective.
The use of a domain-specific image encoder significantly boosts the performance of MAIRA-1. One aspect of this may be the increased image resolution of RAD-DINO (518px), enabling the detection of potentially small and/or subtle findings such as pneumothorax. A larger set of image tokens enabled us to fruitfully scale up the adapter layer, raising the prospect of further gains from yet more complex processing of image tokens.
By using GPT-3.5 to paraphrase reports in MIMIC-CXR, we observe a boost in clinical metrics with a minor degradation of lexical metrics. We speculate that this paraphrasing serves as a semantics-preserving transformation of the text, encouraging the model to focus on the key aspects of the report without overfitting to its style.
Our stratified analysis reveals significantly higher performance on studies containing an Indication section. We hypothesise two mechanisms for this. Firstly, knowledge of the ‘question’ behind the report may inform the expected Findings; for example, ‘…please evaluate for pleural effusion’ (Indication) prompts: ‘Left mid to lower lung opacified is likely a combination of moderate right-sided pleural effusion…’ (Findings). Secondly, the Indication section can include additional clinical context on the patient, which has been shown to improve interpretation (Yapp et al., 2022).
By stratifying by findings class, we observe that the behaviour of MAIRA-1 varies, yielding the best overall performance on classes such as support devices, pleural effusion, and cardiomegaly. For certain clinically actionable findings such as edema and consolidation, MAIRA-1 exhibits unsatisfactory recall, albeit with consistently high negative predictive value. This highlights that reporting aggregate metrics alone may obscure disparate performance within findings classes or patient subgroups.
To understand whether models such as MAIRA-1 are clinically useful, more fine-grained metrics, categories, and exemplar datasets are important, as well as evaluations in realistic use-contexts (Huang et al., 2023). For example, the commonly used CheXpert classes include ‘Pneumonia’ despite it being a clinical diagnosis which should not be assessed from an image alone. Efforts such as RadCliQ are valuable towards the development of radiology-specific evaluation methods, but, as highlighted in our error analysis, the presence of incomplete or otherwise ‘imperfect’ reports in the dataset mean correctly generated reports may still be unfairly penalised.
Reports often contain information derived from prior studies, clinical notes, accompanying laterals or other relevant imaging examinations of the patient. Models such as MAIRA-1 and much prior work, trained to generate reports from single images, are thus forced to hallucinate this information from their limited context. Existing commonly reported metrics do not directly quantify this failure mode. Future versions of MAIRA-1 could include the current and previous study, thereby reducing the need to hallucinate, as demonstrated in Bannur et al. (2023a).
7 Conclusion
We have presented MAIRA-1 as a proof-of-concept for a radiology-adapted large multimodal model. Despite training on a relatively small dataset with a conventional language-modelling loss, it exhibits competitive performance with existing state-of-the-art on findings generation across a broad suite of metrics, benefiting from a domain-specific image encoder, simple training paradigm and text-based augmentation. We believe the performance and clinical utility of MAIRA-1 can be pushed much further by allowing the model to consider multiple images, e.g. priors and complementary views, and by training on larger, more diverse, and higher-quality datasets.
Appendix
Appendix A Do we need to pre-train the adapter?
Prior work conducted an initial training of just the adapter layer, keeping the LLM frozen [Liu et al., 2023b, Li et al., 2023]. In Table 8, we investigate this choice. We start from a baseline model with CLIP-ViT-L-336px as image encoder and Vicuna-7B as LLM. We obtain a fine-tuned adapter by training on the findings generation task. We then compare the impact of using this adapter against a randomly initialised adapter of the same size in our typical training setup. Although prior work often uses this two-stage training process, we find that initializing from the pretrained adapter causes the final performance to be significantly worse.
Category | Metric | Random adapter init | Pretrained adapter init |
Lexical | ROUGE-L | 28.2 [27.8, 28.8] | 27.4 [27.0, 27.9] |
BLEU-1 | 32.8 [32.0, 33.5] | 31.6 [30.9, 32.3] | |
BLEU-4 | 12.7 [12.2, 13.2] | 12.1 [11.7, 12.7] | |
METEOR | 30.3 [29.8, 30.9] | 29.2 [28.6, 29.7] | |
Clinical | RadGraph-F1 | 20.2 [19.7, 20.8] | 18.4 [17.8, 18.9] |
RGER | 25.0 [24.4, 25.6] | 23.0 [22.5, 23.7] | |
CheXbert vector | 39.6 [38.8, 40.4] | 35.5 [34.6, 36.5] | |
RadCliQ () | 3.29 [3.25, 3.32] | 3.42 [3.39, 3.46] | |
Macro-F1-14 | 29.4 [27.7, 30.8] | 24.0 [22.6, 25.4] | |
Micro-F1-14 | 46.4 [45.0, 47.6] | 39.7 [38.4, 41.0] | |
Macro-F1-5 | 40.7 [38.8, 42.6] | 32.6 [30.9, 34.3] | |
Micro-F1-5 | 48.1 [46.6, 49.8] | 40.1 [38.2, 41.9] |
Appendix B Are gains from GPT-augmentation due to training longer?
We see from Table 4 that adding GPT augmentation improves the performance of the model on clinical metrics. Hoewever, adding GPT paraphrased reports also increases the number of samples in the dataset and thus the training steps, as we train for the full 3 epochs even after adding the paraphrased example to our dataset. To examine whether the gains we get from using GPT-augmented data are purely due to training for a larger number of steps, we run an ablation. We create a variant dataset with the same set of samples as that from GPT-augmentation, but using the original report rather than the GPT-paraphrased variant. In Table 9 we compare this setting with the use of GPT-paraphrased reports.
Category | Metric | MAIRA-1 | Control |
Lexical | ROUGE-L | 28.9 [28.4, 29.4] | 28.5 [28.0, 28.9] |
BLEU-4 | 14.2 [13.7, 14.7] | 14.0 [13.6, 14.5] | |
METEOR | 33.3 [32.8, 33.8] | 32.7 [32.2, 33.2] | |
Clinical | RadGraph-F1 | 24.3 [23.7, 24.8] | 22.8 [22.1, 23.3] |
RGER | 29.6 [29.0, 30.2] | 27.8 [27.3, 28.4] | |
CheXbert vector | 44.0 [43.1, 44.9] | 42.8 [42.0, 43.6] | |
RadCliQ () | 3.10 [3.07, 3.14] | 3.16 [3.13, 3.19] | |
Macro-F1-14 | 38.6 [37.1, 40.1] | 36.7 [35.0, 38.2] | |
Micro-F1-14 | 55.7 [54.7, 56.8] | 54.3 [53.3, 55.4] | |
Macro-F1-5 | 47.7 [45.6, 49.5] | 45.2 [43.3, 47.1] | |
Micro-F1-5 | 56.0 [54.5, 57.5] | 54.4 [53.0, 55.8] |